YoursTruly:

Leigh took out a payday loan for $400 due in 2 weeks that charged a $50 fee. What is the periodic interest rate of the loan? Show your work.

YoursTruly:

@Arcadiouse

Arcadiouse:

weird question idk

RawlsK29:

You would start out with the formula.

RawlsK29:

So, the fee is $50 and the P which would be the loan is $400 and R is the periodic interest rate. Now, you have to solve T times 2/50 or T*2/50 You would now solve for the formula R R=I÷pt ( We got the T ready, now we need the R ) So we now have R=50÷(400×(2÷52)) because we had 2/50 Then we multiply R=3.25×100 R=325% What is the periodic interest rate of the loan? 325% (How I got 3.25 is I did 50÷(400×(2÷52) Hope this Helps!(: